OpenStudy (anonymous):

How to solve this integral indefinite Integral of sin(x)/cos^2(x) without using 'u' substitution?

jimthompson5910 (jim_thompson5910):

hint: \[\Large \frac{\sin(x)}{\cos^2(x)} = \frac{\sin(x)}{\cos(x)}*\frac{1}{\cos(x)}\]

OpenStudy (anonymous):

\[\int\limits_{a}^{b}\frac{ \sin x }{ \cos ^2(x) }\]

OpenStudy (anonymous):

Ah it makes sense now

jimthompson5910 (jim_thompson5910):

I'm glad it does